OpenStudy (anonymous):

find the indicated derivatives f(x)=1/2x+3, y''

OpenStudy (anonymous):

Just to be sure, is this f(x) = (1/2)x + 3, or f(x) = 1/(2x +3)?

OpenStudy (anonymous):

its f(x)=1/(2x+3)

OpenStudy (anonymous):

well then, we have \[f'(x) = (-1/(2x + 3)^2)\times 2 = -2/(2x+3)^2\] where the multiplication by 2 occurs by the chain rule and then \[f''(x) = (4/(2x+3)^3)\times 2 = 8/(2x+3)^3\]

OpenStudy (anonymous):

ohh okay! thanks so much!